OpenStudy (anonymous):

∫|2-x|dx intregation limits is 0 to 2?

OpenStudy (anonymous):

\[\int_0^2|2-x|dx=\int_0^2(2-x)dx\]

OpenStudy (anonymous):

\[\int\limits_{0}^{2} (2-x)dx\] \[\int\limits_{0}^{2} 2x- x^2 /2\] \[(2(2)-(2)^2 / 2 ) - ( 2(0) - (0)^2/2)\] 2

OpenStudy (anonymous):

2 is the answer